OpenStudy (anonymous):

Write the equation in slope-intercept form for the following information: b = 4 and m = -1

OpenStudy (anonymous):

y=-x+4 y=mx+b

OpenStudy (anonymous):

y=-1x+4

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

np <3